# OpenFACADES: An Open Framework for Architectural Caption and Attribute Data Enrichment via Street View Imagery

### v1

@@ 2.2 Vision models in urban analytics
With the rapid development of deep learning techniques over the past decade, diverse methods have been developed to extract urban cues from visual information, providing efficient and scalable frameworks for understanding built environments. In terms of building facade research, in particular, Convolutional Neural Networks (CNNs) have been widely employed due to their strong feature representation capabilities. Among them, VGG, DenseNet, and ResNet have been extensively applied to achieve, or serve as benchmarks for, the accurate classification and evaluation of building functions (Kang et al., 2018), materials (Ghione et al., 2022; Raghu et al., 2023), architectural styles (Lindenthal and Johnson, 2021; Sun et al., 2022b; Ogawa et al., 2023), and human perceptions (Liang et al., 2024). Additionally, Vision Transformers (ViTs) have emerged as powerful alternatives, leveraging self-attention mechanisms to capture long-range dependencies in building images. Recent studies have demonstrated the effectiveness of ViTs in urban analytics, achieving state-of-the-art performance in material recognition, and construction period prediction (Raghu et al., 2023; Ogawa et al., 2023). Beyond that, hybrid models combining various model backbones have been further developed to consider multi-dimensional features as input, improving comprehensiveness and generalizability in multi-scale urban analysis (Huang et al., 2023; Jia et al., 2024; Fujiwara et al., 2024).
@@
Rapid advancements in LLMs offer new avenues for extracting nuanced insights about complex urban environments. Notably, VLMs combine visual and linguistic modalities, leveraging deep semantic reasoning to establish rich connections between visual concepts and textual descriptions (Wu et al., 2023a; Li et al., 2024a). Building on these capabilities, recent work in remote sensing demonstrates how VLMs can exceed traditional CV methods by producing more context-aware and human-like interpretations (Al Rahhal et al., 2022; Zia et al., 2022; Hu et al., 2023), thereby providing not only precise visual recognition but also a semantic understanding of objects and their relationships within complex environments. In terms of street-level building research, recent studies have explored the state-of-the-art models for automated building annotation. For example, Li et al. (2024b) employed ChatGPT-4o to generate structured multi-label annotations for buildings using SVIs across multiple cities. Similarly, Zeng et al. (2024) assessed the model’s performance in zero-shot building age prediction, finding that ChatGPT-4 effectively estimates the construction period of buildings based on a single-perspective view. However, deploying proprietary LLMs such as ChatGPT-4o at scale presents limitations. Model inference relies on API-based access, which incurs high computational costs, making large-scale applications financially and computationally restrictive, which also constrains the efficiency for fine-tuning, limiting their adaptability for domain-specific urban studies. To address these challenges, recent open-source initiatives have produced diverse series of LLMs, including Qwen-VL (Wang et al., 2024b), Llama (Dubey et al., 2024), and InternVL (Chen et al., 2024b), enabling greater customization and efficiency in downstream tasks. These models exhibit unified capabilities to process multi-dimensional inputs, generating context-aware descriptions informed by their pretraining on large-scale, diverse datasets. By effectively capturing the distributions of natural language and multimodal semantics, open-source VLMs exhibit strong generalization performance while maintaining the flexibility for task-specific adaptations. This capability holds significant potential for advancing street-level urban analysis, as their ability to interpret human-centric observations closely aligns with how individuals perceive and contextualize the built environment.
@@
Hence, we propose a reproducible methodology for integrating open-source multimodal building data from global cities into a comprehensive dataset, incorporating objective attributes and detailed VLM-based interpretations. Table 1 provides an overview of existing SVI datasets related to building attributes, highlighting how our contribution addresses current limitations while significantly expanding the scale, scope, and dimensionality of SVI-based datasets for building-related research. This advancement not only enhances the accessibility and adaptability of building datasets but also paves the way for broader, more inclusive, and scalable applications in urban analytics.
@@
Lindenthal and Johnson (2021) image classification age GSV perspective 29,177 1 Europe Georgian, early Victorian, late Victorian/Edwardian, interwar, postwar, contemporary, revival (7 categories). Raghu et al. (2023) image classification surface material GSV perspective 985 3 Asia, North America, Europe brick, stucco, rustication, siding, wood, metal, other (7 categories) SVI4BuildingFunc (Li et al., 2025b) object detection type GSV panoramic 15,400 4 North America, Europe varies by city (e.g., high residential, low residential, commercial, office, walk-up buildings, mixed-up buildings; 5 to 6 categories per city) OpenFACADES Image labeling and captioning type, age, floor, surface material, feature description Mapillary individual building images 30,180 7 North America, Europe, Asia Type: apartments, house, retail, office, hotel, industrial, religious, education, public, garage (10 categories);
@@ 3 Methodology
In this study, we introduce OpenFACADES, a comprehensive framework for acquiring building images from Street View Imagery (SVI) and automatically annotating them with crowdsourced data, supported by a series of methodological advancements. This framework facilitates the development of large multimodal models tailored for architectural attributes question-answering and captioning. The framework is structured into three main steps, as illustrated in Figure 1:
(1) Integrating multimodal crowdsourced data. Initially, crowdsourced SVI metadata and building data are collected for the designated research areas. Then, isovist analysis is performed to simulate the theoretical angles of view (AOV) from each camera location to the target structures. Based on observation quality, high-quality SVIs are retrieved and further filtered based on their image features, ensuring that only candidate images with optimal visibility and clarity are retained for subsequent analysis.
(2) Retrieving building image data. Based on the geospatial AVOs simulated, we map the relative viewing angles and detect target buildings within the image space. This process enable us to precisely identify associate building information with their visual representations. Then, based on the coordinates of bounding boxes, building images are reprojected from panoramic to perspective view, generating observation aligned with real world observation. These images further undergo a image filtering process to identify high-quality and suitable building images.
(3) Establishing dataset and multimodal models. Building images with available crowdsourced data form a dataset with four label types: categorical, single-word Q&A, multi-attribute Q&A, and captioning. Categorical labels are derived from building information, while single-word Q&A labels append categorical labels to targeted questions, generating concise question-to-label pairs. Multi-attribute Q&A and captioning labels are generated using ChatGPT-4o, enabling detailed textual descriptions and structured annotations for comprehensive building attribute analysis. The last three label types are utilized to fine-tune vision-language models, enabling a versatile model for multi-attribute building labeling and captioning with enhanced contextual understanding.
@@ 3.1 Integrating multimodal crowdsourced data
The workflow of integrating multimodal crowdsourced data for building analysis is illustrated in Figure 2. The process includes: (1) preprocessing street-level imagery based on metadata; (2) refining building footprints and attributes; (3) calculating angle of view (AOV) to assess building visibility; and (4) selecting high-quality target images based on AOV thresholds and quality metrics.
Image data preparation
At the first stage, the raw metadata of street-level image data from crowdsourced platform is obtained within study areas before requesting the images. Here, Mapillary is chosen for its extensive global coverage, high-quality user-generated content, and open-access policies that enable reproducible and scalable urban research (Hou and Biljecki, 2022; Kapp et al., 2025; Danish et al., 2025). Specifically, the metadata — comprising location coordinates (computed_geometry), compass angle (computed_compass_angle), capture time (captured_at), and quality indicator (quality_score) — is utilized to structure sorting and quality assessments. Here, filtering operations remove images captured outside the defined study area, exclude multiple images from the same spatial point to prevent redundant viewpoints, and discard those with poor resolution or quality defects. The output of this phase is a curated set of image metadata, with their corresponding unique image IDs, coordinates, and compass angles, prepared for subsequent spatial analyses.
@@
Building data preparation
@@
In parallel, building geometries and associated metadata, retrieved from OpenStreetMap (OSM) [^1] https://openstreetmap.org/, are selected as the base dataset for the footprint layer and building information to undergo a refinement process. Based on the context of cities, data harmonization is conducted to supplement missing building footprints from OSM, as well as supplement insufficient building attributes from other data sources, such as Overture Maps [^2] https://overturemaps.org/ and government datasets. Attributes commonly include unique identifiers, building type, facade material, number of floors, construction dates, and polygon geometries. Inconsistencies and outliers — such as footprints representing insignificant or extraneous structures (e.g., roof and underground structures), duplicates introduced by overlapping contributions, or buildings located outside the target region — are systematically removed. After applying these filters, the remaining dataset delivers a precise, consistent, and high-quality representation of the built environment, ready for geometric calculations and alignment with the image data.
Isovist analysis
@@
With both image and building datasets prepared, the next step involves performing isovist analysis to compute theoretical AOV from each camera location to the target structures, building on previous studies (Lindenthal and Johnson, 2021; Ogawa et al., 2023; Fan et al., 2025). This analysis identifies each building’s perimeter segments that fall within the camera’s potential field of view and evaluates the observation efficiency of buildings from specific vantage points. First, a search radius of 50 meters is established to identify surrounding buildings from the SVI capture points. Second, sampling points are generated along the polygonal geometries of buildings within the distance threshold, and lines of sight are constructed towards all sampled points of the target buildings. Third, lines of sight intersecting with surrounding building footprints are filtered out, leaving only the largest angular span between the unobstructed lines, which represents the AOV to a building from a given image shooting point. Additionally, the left and right boundaries of the AOV are recorded as azimuth angles relative to the true north, providing detailed spatial orientation for subsequent tasks. This process identifies which buildings are potentially visible from each image capture point, thereby aligning the building information with the corresponding imagery metadata.
Candidate image selection
@@
Based on the theoretical visibility of buildings, the final stage identifies candidate images most likely to provide reliable and interpretable observations. Criteria derived from the absolute AOV eliminate images taken at excessive distances, those with extreme observation angles (in this study, AOV greater than 120 degrees or smaller than 10 degrees are filtered out), and images that essentially duplicate prior perspectives. Given that crowdsourced SVI, as typical for volunteered geographic information (VGI), can vary in quality and may contain various errors (Hou and Biljecki, 2022), the selected images are then retrieved from Mapillary and evaluated against additional quality metrics — such as brightness, sharpness, and visual complexity — to further identify suitable images for the dataset. Images captured under unsuitable conditions (e.g., nighttime, severe overexposure) or containing excessive visual clutter are removed based on the CV models released in NUS Global Streetscapes (Hou et al., 2024). The result is a high-quality, focused selection of candidate street-level images, optimized for integration with building data in subsequent object detection workflows.
@@ 3.2 Retrieving building image data
Figure 3 demonstrates the pipeline for extracting and selecting building images from street-level imagery. The process consists of three main steps: Building detection, image reprojection and image filtering:
Building detection
@@
First, we interpret the bounding box region in terms of pixel coordinates within the panoramic imagery, obtaining the box center as ($c_{u}$, $c_{v}$), along with its $width$ and $height$, which are essential for subsequent tasks. Second, a virtual pinhole camera model is constructed based on the specified AOV to a target building and the bounding box $width$. The focal length $f$ and principal point ($c_{x}$, $c_{y}$) in camera coordinate are computed as:
@@
where $c_{u}$ and $c_{v}$ are normalized to a range of [0, 1], with $c_{u}$ as the horizontal center and $c_{v}$ as the vertical center of the bounding box region. The yaw angle $\theta$ defines the horizontal rotation of the camera and spans from $-180^{\circ}$ to $180^{\circ}$. The pitch angle $\phi$ defines the vertical rotation of the camera and spans from $-90^{\circ}$ to $90^{\circ}$. The combined rotation matrix $R$ is formed as the product of two individual rotation matrices based on Rodrigues’ formula: $R_{y}$($\theta$), which rotates the coordinate system around the y-axis (yaw), and $R_{x}$($\phi$), which rotates the coordinate system around the x-axis (pitch). $\mathbf{v}_{cam}$ is the original direction vector in the camera’s coordinate system, while $\mathbf{v}_{rot}$ is the new direction vector after applying the rotations, pointing toward the desired region of the spherical panorama.
@@
At these coordinates, pixel values are sampled from the original panoramic image, and reprojected to generate the rectified perspective view using the remap function from OpenCV library. This transformation eliminates the spherical distortions inherent in panoramic imagery, producing a visually intuitive and geometrically corrected view aligned with the detected object. As examples demonstrated in Figure 4, this correction is crucial not only for preserving essential structural details for model interpretation but also for mitigating distortions that could otherwise misalign architectural features. This preprocessing step enhances the model’s ability to accurately analyze building attributes in urban applications.
@@
Image filtering
The features of the detected individual building images are further analyzed to refine and enhance the image dataset. ZenSVI (Ito et al., 2024), an open-source library for street-level imagery analysis, is integrated into the framework to facilitate the extraction of image features. We analyze image features across four key dimensions: blurriness, brightness, semantic segmentation, and scene classification. These dimensions are utilized to identify high-quality and suitable building images for inclusion in the dataset: blurriness, brightness, semantic segmentation, and scene classification. Blurriness is evaluated using the OpenCV Laplacian operator to filter out images with motion blur or poor focus, while brightness assessment removes those with suboptimal illumination. A pre-trained Place365 model (Zhou et al., 2017) excludes indoor scenes, and semantic segmentation is applied to detect and minimize occlusions (e.g., trees, vehicles, walls), ensuring that selected images predominantly showcase building facades and maintain high visual quality.

### v2

@@ 2.2 Vision models in urban analytics
With the rapid development of deep learning techniques over the past decade, diverse methods have been developed to extract urban cues from visual information. In terms of building facade research, in particular, CNNs have been widely employed due to their strong feature representation capabilities. Among them, VGG, DenseNet, and ResNet have been extensively applied to achieve, or serve as benchmarks for, the accurate classification and evaluation of building functions (Kang et al., 2018), materials (Ghione et al., 2022; Raghu et al., 2023), architectural styles (Lindenthal and Johnson, 2021; Sun et al., 2022b; Ogawa et al., 2023), and human perceptions (Liang et al., 2024). Additionally, Vision Transformers (ViTs) have emerged as powerful alternatives, leveraging self-attention mechanisms to capture long-range dependencies in building images. Recent studies have demonstrated the effectiveness of ViTs in urban analytics, achieving state-of-the-art performance in material recognition, and construction period prediction (Raghu et al., 2023; Ogawa et al., 2023). Beyond that, hybrid models combining various model backbones have been further developed to consider multi-dimensional features as input, improving comprehensiveness and generalizability in multi-scale urban analysis (Huang et al., 2023; Jia et al., 2024; Fujiwara et al., 2024).
@@
Rapid advancements in LLMs offer new avenues for extracting nuanced insights about complex urban environments. Notably, VLMs combine visual and linguistic modalities, leveraging deep semantic reasoning to establish rich connections between visual concepts and textual descriptions (Wu et al., 2023a; Li et al., 2024a). Building on these capabilities, recent work in remote sensing demonstrates how VLMs can exceed traditional CV methods by producing more context-aware and human-like interpretations (Al Rahhal et al., 2022; Zia et al., 2022; Hu et al., 2025), thereby providing not only precise visual recognition but also a semantic understanding of objects and their relationships within complex environments. In terms of street-level building research, recent studies have explored the state-of-the-art models for automated building annotation. For example, Li et al. (2024b) employed ChatGPT-4o to generate structured multi-label annotations for buildings using SVIs across multiple cities. Similarly, Zeng et al. (2024) assessed the model’s performance in zero-shot building age prediction, finding that ChatGPT-4 effectively estimates the construction period of buildings. However, deploying proprietary LLMs such as ChatGPT-4o at scale presents limitations. Model inference relies on API-based access, which incurs high computational costs, making large-scale applications financially and computationally restrictive, which also constrains the efficiency for fine-tuning, limiting their adaptability for domain-specific urban studies. To address these challenges, recent open-source initiatives have produced diverse series of LLMs, including Qwen-VL (Wang et al., 2024b), Llama (Dubey et al., 2024), and InternVL (Chen et al., 2024b), enabling greater customization and efficiency in downstream tasks. These models exhibit unified capabilities to process multi-dimensional inputs, generating context-aware descriptions informed by their pretraining on large-scale, diverse datasets. This capability holds significant potential for advancing street-level urban analysis, as their ability to interpret human-centric observations closely aligns with how individuals perceive and contextualize the built environment.
@@
Hence, we propose a reproducible methodology for integrating open-source multimodal building data from global cities into a comprehensive dataset, incorporating objective attributes and detailed captions. Table 1 provides an overview of existing SVI datasets related to building attributes, highlighting how our contribution addresses current limitations while significantly expanding the scale, scope, and dimensionality of SVI-based datasets for building-related research. This advancement not only enhances the accessibility and adaptability of building datasets but also paves the way for broader, more inclusive, and scalable applications in urban analytics.
@@
Lindenthal and Johnson (2021) image classification age GSV perspective 29,177 1 Europe Georgian, early Victorian, late Victorian/Edwardian, interwar, postwar, contemporary, revival (7 categories). Raghu et al. (2023) image classification surface material GSV perspective 985 3 Asia, North America, Europe brick, stucco, rustication, siding, wood, metal, other (7 categories) SVI4BuildingFunc (Li et al., 2025b) object detection type GSV panoramic 15,400 4 North America, Europe varies by city (e.g., high residential, low residential, commercial, office, walk-up buildings, mixed-up buildings; 5 to 6 categories per city) OpenFACADES Image labeling and captioning type, age, floor, surface material, feature description Mapillary individual building images 31,180 7 North America, Europe, Asia Type: apartments, house, retail, office, hotel, industrial, religious, education, public, garage (10 categories);
@@ 3 Methodology
In this study, we introduce OpenFACADES, a comprehensive framework for acquiring building images from SVI and automatically annotating them with crowdsourced data. This framework facilitates the development of large multimodal models tailored for architectural attributes question-answering and captioning. The framework is structured into three main steps, as illustrated in Figure 1:
(1) Integrating multimodal crowdsourced data. Initially, crowdsourced SVI metadata and building data are collected for research areas. Then, isovist analysis is performed to simulate the theoretical angles of view (AOV) from each camera location to the target structures. SVIs with optimal visibility are then retrieved and filtered by image features to ensure that only high-quality candidates are retained for subsequent analysis.
(2) Retrieving building image data. Based on the geospatial AOVs simulated, we map the relative viewing angles and detect target buildings within the image space. This process enables us to precisely identify associate building information with their visual representations. Then, based on the coordinates of bounding boxes, building images are reprojected from panoramic to perspective view, generating holistic building images. These images further undergo a filtering process to identify high-quality and suitable building views.
(3) Establishing dataset and multimodal models. Building images with available crowdsourced data form a dataset with four label types: single-attribute label, single-attribute Q&A, multi-attribute Q&A, and captioning. Single-attribute labels are derived from building information, while single-attribute Q&A append those attribute labels to targeted questions, generating concise question-to-label pairs. Multi-attribute Q&A and captioning labels are generated using ChatGPT-4o, enabling detailed textual descriptions and structured annotations for comprehensive building attribute analysis. The last three label types are utilized to fine-tune vision-language models, enabling a versatile model for multi-attribute building labeling and captioning with enhanced contextual understanding.
@@ 3.1 Integrating multimodal crowdsourced data
The workflow of integrating multimodal crowdsourced data for building analysis is illustrated in Figure 2. The process includes:
Image data preparation.
At the first stage, the raw metadata of street-level image data from crowdsourced platform is obtained within study areas before requesting the images. Here, Mapillary is chosen for its extensive global coverage, high-quality user-generated content, and open-access policies that enable reproducible and scalable urban research (Hou and Biljecki, 2022; Kapp et al., 2025; Danish et al., 2025). Specifically, the metadata, comprising image type (is_pano), location coordinates (computed_geometry), compass angle
(computed_compass_angle), capture time (captured_at), and quality indicator (quality_score), is utilized to structure sorting and quality assessments. Filtering operations select panoramic imagery (is_pano=True), remove images captured outside the defined study area, exclude multiple images from the same spatial point to prevent redundant viewpoints, and discard those with poor resolution or quality defects. The output of this phase is a curated set of image metadata, with their corresponding unique image IDs, coordinates, and compass angles, prepared for subsequent spatial analyses.
@@
Building data preparation.
@@
In parallel, building geometries and associated metadata are retrieved from OpenStreetMap (OSM) [^1] https://openstreetmap.org/. Data harmonization is then conducted to supplement missing building footprints and insufficient building attributes from other data sources, such as Overture Maps [^2] https://overturemaps.org/ and government datasets. Attributes include unique identifiers, building type, facade material, number of floors, construction dates, and polygon geometries. Inconsistencies and outliers, such as footprints representing insignificant or extraneous structures (e.g., roof and underground structures), duplicates introduced by overlapping contributions, or buildings located outside the target region, are systematically removed. The remaining dataset delivers a precise, consistent, and high-quality representation of the built environment, ready for geometric calculations and alignment with the image data.
Isovist analysis.
@@
With both image and building datasets prepared, isovist analysis is applied to compute theoretical AOV from each camera location to the target structures, building on previous studies (Lindenthal and Johnson, 2021; Ogawa et al., 2023; Fan et al., 2025). This analysis identifies each building’s perimeter segments that fall within the camera’s potential field of view and evaluates the observation efficiency of buildings from specific vantage points. First, a search radius (e.g., 50 meter) is established to identify surrounding buildings from the SVI capture points. Second, sampling points are generated along the polygonal geometries of buildings within the distance threshold, and lines of sight are constructed towards all sampled points of the target buildings. Third, lines of sight intersecting with surrounding building footprints are filtered out, leaving only the largest angular span between the unobstructed lines, which represents the AOV to a building from a given image shooting point. Additionally, the left and right boundaries of the AOV are recorded as azimuth angles relative to the true north, providing detailed spatial orientation for subsequent tasks. This process identifies which buildings are potentially visible from each image capture point, thereby aligning the building information with the corresponding imagery metadata.
Candidate image selection.
@@
Based on the theoretical visibility of buildings, the final stage identifies candidate images most likely to provide reliable observations. Criteria derived from the absolute AOV values are used to eliminate images taken from excessive distances, extreme viewing angles, or redundant perspectives. The remained imagery IDs are then used to retrieve image data from Mapillary. Given that crowdsourced SVI can vary in quality and may contain various errors (Hou and Biljecki, 2022), these images are further assessed using quality metrics such as brightness, sharpness, and visual complexity to determine their suitability for inclusion. Images captured under unsuitable conditions (e.g., nighttime, severe overexposure) or containing excessive visual clutter are removed based on the CV models released in NUS Global Streetscapes (Hou et al., 2024). The result is a curated set of candidate street-level images, optimized for integration with building data in subsequent object detection workflows.
@@ 3.2 Retrieving building image data
Figure 3 demonstrates the pipeline for extracting and selecting building images from street-level imagery. The process consists of three main steps: Building detection, image reprojection and image filtering.
Building detection.
@@
First, we interpret the bounding box region in terms of pixel coordinates within the panoramic imagery, obtaining the box center as ($c_{u}$, $c_{v}$), along with its widthwidth and heightheight, which are essential for subsequent tasks. Second, a virtual pinhole camera model is constructed based on the specified AOV to a target building and the bounding box widthwidth. The focal length $f$ and principal point ($c_{x}$, $c_{y}$) in camera coordinate are computed as:
@@
where $c_{u}$ and $c_{v}$ are normalized to a range of [0, 1], with $c_{u}$ as the horizontal center and $c_{v}$ as the vertical center of the bounding box region. The yaw angle $\theta$ defines the horizontal rotation of the camera and spans from $180^{\circ}$ to $180^{\circ}$. The pitch angle $\phi$ defines the vertical rotation of the camera and spans from $90^{\circ}$ to $90^{\circ}$. The combined rotation matrix $R$ is formed as the product of two individual rotation matrices based on Rodrigues’ formula: $R_{y}$($\theta$), which rotates the coordinate system around the y-axis (yaw), and $R_{x}$($\phi$), which rotates the coordinate system around the x-axis (pitch). $\mathbf{v}_{cam}$ is the original direction vector in the camera’s coordinate system, while $\mathbf{v}_{rot}$ is the new direction vector after applying the rotations, pointing toward the desired region of the spherical panorama.
@@
At these coordinates, pixel values are sampled from the original panoramic image, and reprojected to generate the rectified perspective view using the remap function from OpenCV library. This transformation eliminates the spherical distortions inherent in panoramic imagery, producing a visually intuitive and geometrically corrected view aligned with the detected object. As examples demonstrated in Figure 4, this correction is crucial not only for preserving essential structural details for model interpretation but also for mitigating distortions that could otherwise misalign architectural features. This preprocessing step enhances the model’s ability to accurately analyze building attributes.
@@
Image filtering.
The features of the detected individual building images are further analyzed to refine and enhance the image dataset. ZenSVI (Ito et al., 2025), an open-source library for street-level imagery analysis, is integrated into the framework to facilitate the extraction of image features. We analyze image features across four key dimensions: blurriness, brightness, semantic segmentation, and scene classification. These dimensions are utilized to identify high-quality and suitable building images for inclusion in the dataset. Blurriness is evaluated using the OpenCV Laplacian operator to filter out images with motion blur or poor focus, while brightness assessment removes those with suboptimal illumination. A pre-trained Place365 model (Zhou et al., 2017) excludes indoor scenes, and semantic segmentation is applied to detect and minimize occlusions (e.g., trees, vehicles, walls), ensuring that selected images predominantly showcase building facades and maintain high visual quality.